I got solutions but wanted to check!
Find all the real solutions of the following equation by first rewriting the equation as a quadratic equation.
please explain your steps if your would?
s^4 - 26s^2 + 25 = 0
thank you for your time!
my answers were radical 1 and -radical 5
cz

Answer by mananth(16946)

s^4 - 26s^2 + 25 = 0
s^4-25s^2-s^2+25=0
s^2(s^2-25)-1(s^2-25)=0
(s^2-25)(s^2-1)=0
(s+5)(s-5)(s+1)(s-1)=0
..
s= + / - 5 , +/- 1